OpenStudy (narusamisty):

Please Help will medal!!! :D 1. The scores for all high school seniors taking the verbal section of the Scholastic Aptitude Test (SAT) is a normal distribution and in a particular year had a mean of 490 and a standard deviation of 100. a.If a student had a z-score of 1.8, how much higher would he have scored on the test than someone who had a z-score of -.3?

jimthompson5910 (jim_thompson5910):

`.If a student had a z-score of 1.8` then what is the raw score x? Use this formula to determine the answer \[\Large z = \frac{x - \mu}{\sigma}\] In this case, \[\Large \mu = \text{mean} = 490\] \[\Large \sigma = \text{standard deviation} = 100\]

jimthompson5910 (jim_thompson5910):

\[\Large z = \frac{x - \mu}{\sigma}\] \[\Large 1.8 = \frac{x - 490}{100}\] solve for x

OpenStudy (narusamisty):

Would the raw score be -4.9? :)

jimthompson5910 (jim_thompson5910):

no, it should be close to 490

jimthompson5910 (jim_thompson5910):

the positive z score means the raw score (x) is above 490

jimthompson5910 (jim_thompson5910):

Step 1) z score formula Step 2) plug in the given values Step 3) Multiply both sides by 100 Step 4) Multiply and simplify (notice the cancelation on the right side) Step 5) Add 490 to both sides Step 6) Combine like terms \[\Large {\color{green}{\text{Step 1)}}} \ \ z = \frac{x - \mu}{\sigma}\] \[\Large {\color{green}{\text{Step 2)}}} \ \ 1.8 = \frac{x - 490}{100}\] \[\Large {\color{green}{\text{Step 3)}}} \ \ 100*1.8 = 100*\frac{x - 490}{100}\] \[\Large {\color{green}{\text{Step 4)}}} \ \ 180 = x-490\] \[\Large {\color{green}{\text{Step 5)}}} \ \ 180+490 = x-490+490\] \[\Large {\color{green}{\text{Step 6)}}} \ \ 670 = x\] So the z score of z = 1.8 corresponds to the raw score of x = 670

OpenStudy (narusamisty):

ohh ok i was looking for the raw score of z

jimthompson5910 (jim_thompson5910):

If the z-score is -0.3, then what is the raw score x?

jimthompson5910 (jim_thompson5910):

z = -0.3 means x = ???

OpenStudy (narusamisty):

460 right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

z = -0.3 leads to x = 460 notice how this raw score is below the mean (490)

OpenStudy (narusamisty):

yes

jimthompson5910 (jim_thompson5910):

so you'll now subtract the raw scores x = 670 and x = 460

OpenStudy (narusamisty):

210? :)

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

210 is your final answer

OpenStudy (narusamisty):

ah ok thankyou! is it ok if i ask for help on another question :)

jimthompson5910 (jim_thompson5910):

sure go ahead

OpenStudy (narusamisty):

2. Bob takes a test in English and gets an 85%. The mean of the English class is 76% with a standard deviation of 5%. Sally takes a math test and she gets an 85%. The mean in math class is 62% with a standard deviation of 12%. Who did relatively better on their test compared to their classmates? Explain your answer using z-scores.

jimthompson5910 (jim_thompson5910):

You'll use this formula again \[\Large z = \frac{x - \mu}{\sigma}\]

jimthompson5910 (jim_thompson5910):

The english class has mu = 0.76 sigma = 0.05 what is the z score if the raw score is x = 0.85 ?

OpenStudy (narusamisty):

1.8? :)

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

The math class has mu = 0.62 sigma = 0.12 if x = 0.85 then z = ???

OpenStudy (narusamisty):

1.92 right ?

jimthompson5910 (jim_thompson5910):

yes after you round

jimthompson5910 (jim_thompson5910):

so because Sally has the higher z-score, she did better relative to her classmates

OpenStudy (narusamisty):

Thankyou so much!!

jimthompson5910 (jim_thompson5910):

no problem